OpenStudy (anonymous):

When the following quadratic equation is written in general form, what is the value of "c"?

OpenStudy (anonymous):

\[-\frac{ 3 }{ 4 }x^{2}+2=0\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@iGreen is -2 the answer?

OpenStudy (anonymous):

the answer is 2 :)

OpenStudy (anonymous):

-8 -6 -2 those are my choices @jo.harvey

OpenStudy (anonymous):

-8 is correct too yes

OpenStudy (anonymous):

than a becomes 3

OpenStudy (anonymous):

ur confusing me

OpenStudy (anonymous):

now a = -3/4 and c=2 if than you multiply the entire equation by -4 you get a=3 and c=-8